Divide:
x^2 + 12x + 35/x^2 + 4x - 5 / x^2 + 3x - 28 / x-1
Find (f + g)(7) when f(x)= x + 5 and g(x)= -x^2 - 6


I tried wroking out the first problem and kept getting (x+5)/(x+5)(x+4)
Im not sure if the x+5 cancels and leaves x+4. Please help. And for the second problem im confused on what to do. Please help.

Answer by stanbon(75887)

Factor where you can to get:
[(x+7)(x+5)]/[(x+5)(x-1)]/[(x+7)(x-4)]/(x-1)
This is like a/b/c/d = (a/b)*(d/c)
So, cancel factors that are common to a numerator and a denominator,
for example cancel (x+5) and (x-1) and (x+7) to get:
1/(x-4)
(f + g)(7)= f(7) + g(7)= (7+5) + (-(7^2)-6)
=12-55=-43
Cheers,
Stan H.